In the following questions, give all your answers to two decimals.


Patrice works as an economist for the Bureau of Labor Statistics (BLS). Her current project is to estimate the effect of changes in income, prices of related goods, and the price of potatoes on the demand for beef. Patrice has the following data:

 Price elasticity of demand for beef–0.80Income elasticity of demand for beef+ 1.40Cross-price elasticity between beef and chicken+1.20Cross-price elasticity between beef and potatoes–0.50




1
Expert's answer
2021-03-09T07:39:25-0500

Price elasticity =change in quantity supplied /change in price.

Income elasticity=change in quantity change in income.

Cross price elasticity=change in quantity of good A/change in price of good B


Suppose the price of beef falls by 5%. All else equal, the quantity of beef demanded would fall

  by  6.40  %. 

"8*0.8=6.4"

If consumer income falls by 9%, the quantity of beef demanded would rise

  by 7.00   %.

"1.4*9=7"

If the price of potatoes rises by 20%, the quantity of beef demanded would  fall

  by   10.00 %.

"20*0.5=10"